OpenStudy (anonymous):

which of the following lines is perpendicular to the line y=2/3 x+6?

OpenStudy (anonymous):

Choices?

OpenStudy (anonymous):

which liens?

OpenStudy (anonymous):

Even without choices, you should learn the general rule for finding out what a perpendicular line would look like.

OpenStudy (anonymous):

1.) y=2/3 x-2 2.) y= -3/2x-4 3.)x= 2/3 y=1 4.) y= -2/3 x+3

OpenStudy (anonymous):

What is the slope \( y=(2/3) x+6\)?

OpenStudy (anonymous):

oops 3 is x= 2/3 y+1

OpenStudy (anonymous):

A perpendicular line's slope is going to be the negative reciprocal.

OpenStudy (anonymous):

whats the answer?